OpenStudy (anonymous):

y = x² + 3x – 4 y = 2x + 2 (–3, 6) and (2, –4) (–3, –4) and (2, 6) (–3, –4) and (–2, –2) no solution

OpenStudy (campbell_st):

equate the two equations 2x + 2 = x^2 +3x - 4 then x^2 +x -6 = 0 solve by factorising (x - 3)(x +2)=0 then x = 3 or -2

OpenStudy (campbell_st):

substitute each x value to find the corresponding y

hero (hero):

None of them are matching up...unless a mistake was made.

OpenStudy (dumbcow):

x^2 +x -6 = 0 (x+3)(x-2) = 0 x = -3,2

hero (hero):

Another goof up by me